Two inertial systems are uniformly separating at a speed of exactly √0.84𝑐. In one system a jogger runs a mile (1609m) 

in 6 min along the axis of relative motion. How far in meters does he run and how long does it take, to observers in the other 

system?


1
Expert's answer
2021-04-22T07:34:35-0400

To be given in question

Velocity (v)=√(0.84c)

"L_{0}=1609 meter"

"Z_{0}" = 6 minutes

To be asked in question

Distance run=?

Times run=?

We know that

"L=L_{0}\sqrt{1-\frac{v^2}{c^2}}\rightarrow(1)"

"Z=\frac{Z_{0}}{\sqrt{1-\frac{v^2}{c^2}}}\rightarrow(2)"

Eqution (1) put values"L=1609\times\sqrt{1-\frac{\sqrt{.84c}^2}{c^2}}"

"L=1609\times\sqrt{.16}"

"L=1609\times0.4"

"L=643.6meter"

Time

Eqution (2)put values

"Z= \frac{360}{\sqrt{1-\frac{v^2}{c^2}}}"

Put v=√(.84c)

"Z=\frac{360}{0.4}"

Z=900 sec

"Z=\frac{900}{60}" =15 minutes

Z=15minutes after and distance L=643.6meter after observer is other systems "\u2206d=1609-643.9"

Separation of speed =

"\frac{distance}{time}\rightarrow equation (1)"


Put values

"Speed=\frac{965.1}{900}=1.07mete/sec"

"Speed" separation of distance "∆d=d_{1}-d_{2}\rightarrow(2)"

Put valuesd1=1609meter

d2=643.6meter equation (2) put values

∆d=965.1meter

Separation distance 965.1meter

Separation Speed=1.072meter/sec